### v2

# P.O.Box 19395-5531, Tehran, Iran

Abstract
Although previous results have ruled out the possibility of a static horizon in cosmology, we present black hole and white hole metrics that retain static horizons while reproducing cosmological behavior at large distances. Using an appropriate coordinate choice, we demonstrate that a static horizon can exist in a cosmological setting without introducing curvature invariant singularities at the horizon. The resulting metric reduces to the Schwarzschild–de Sitter solution when the Hubble parameter is constant. We find that white hole metrics in an expanding universe (or black holes in a contracting universe) are significantly easier to construct, as a black hole in an expanding cosmology requires the velocity function to change sign. Consequently, this work primarily examines white holes in expanding cosmologies as a foundation for subsequent analysis of black holes in expanding universes. In later sections, we investigate scenarios involving a white hole coupled with cosmological matter, as well as a white hole with both matter and a cosmological constant. Assuming the pressure component takes its cosmological value, we show that the physical radius of the apparent horizon can asymptotically approach a constant value at late times. This metric avoids pathologies such as a singular horizon in the limit of a vanishing Hubble parameter. Finally, we analyze the realistic case of a black hole embedded in pressure-less cosmological matter with and without a cosmological constant and explore its properties.
1 Introduction
Black holes are an indispensable part of any textbook on general relativity, which has established the fundamental mathematical tools in our understanding of the universe as a whole or in other words cosmology, and yet a consistent black hole solution (with or without a static horizon) that can be embedded in cosmology is still a matter of debate (see for instance McVittie ; Einstein:1945id ; Kaloper ; Davidson ; Faraoni ; Gaur ; Dahal ; Ion ; Poplawski ; Farrah ; Cadoni1 ; Cadoni2 ). Indeed there are arguments that forbid the existence of black holes with static horizons, independent of the underlying gravitational theory Davidson Faraoni . Our main initiative in this work is to propose a way out of these no-go results and show that it is possible to find such metrics that behave like a black hole close to the horizon and tend to standard cosmology away from it, meanwhile keeping the horizon static. To this end, in section II, we use the Painlevé-Gullstrand coordinates for both the black hole and cosmological metrics. Due to the simplicity of deriving a white hole in expanding cosmology (or similarly, a black hole in contracting cosmology), we first consider this case and study the possibility of a static horizon. In order to achieve a static horizon, we drop one of the assumptions that led to McVittie's metric. We let $p_{\perp}$ be different from $p_{r}$ for radial distance $r$ small compared to the cosmological horizon (i.e., in the vicinity of the white hole). Assuming our gravitational theory is general relativity, energy density everywhere is the same as the cosmological value, and the other components also tend to cosmological values for $\frac{l_{1}}{r}\ll 1$, where $l_{1}$ is the static white hole horizon radius. This metric satisfies the null convergence condition, i.e. null energy condition (NEC) outside the horizon, but it violates it inside the white hole horizon. To resolve this issue, we drop another one of our assumptions, namely, we allow the energy density to be different from the FLRW one, and show that it is, in principle, possible to have a metric that satisfies NEC inside the white hole horizon as well. We then consider the case of a black hole in expanding cosmology that requires a sign change in the velocity function $h$ by considering a more generic metric ansatz.
In the second part of this work (section III), we consider a more realistic setup with a constant equation of state. Specializing to the case of white hole plus cosmic matter in an expanding cosmology, we find exact white hole solutions and show that it is possible to have a cosmological white hole metric which has an asymptotically (in $t$) constant physical radius. We repeat this calculation for the case of white hole plus matter, plus cosmological constant, and find similar results. In particular, we show that the apparent horizons in the latter case tend to the static horizons of the Schwarzschild-de Sitter metric for large $t$. We then consider a more generic metric ansatz that includes a curvature-like behavior (relative to a flat FLRW metric) and find a black hole metric in matter-dominated expanding cosmology that has a sign change in the velocity function $h$. We conclude in Section IV.
2 The feasibility of a static horizon
In this section, we show that it is in principle possible to have a static horizon in cosmology.
2.1 A naive extension of the Schwarzschild-de Sitter metric
The Schwarzschild metric can be written in Painlevé-Gullstrand coordinates as Painleve Gullstrand
$ds^{2}=-(1-\frac{l_{s}}{r})dt^{2}+2\sqrt{\frac{l_{s}}{r}}drdt+dr^{2}+r^{2}d\Omega_{2}^{2}.$ (1)
On the other hand, the cosmological FLRW metric can be written in Painlevé-Gullstrand coordinates as Gaur2
$ds^{2}=-(1-\frac{\dot{a}^{2}}{a^{2}}r^{2})dt^{2}-2\frac{\dot{a}}{a}rdtdr+dr^{2}+r^{2}d\Omega_{2}^{2}.$ (2)
where $a(t)$ is the scale factor. This coordinate system is specially suitable for writing a metric that interpolates between the cosmological metric and the Schwarzschild metric due to the time-independent radial factor of angular coordinates.
Abiding by the requirement that any cosmological black hole metric should tend to the Schwarzschild-de Sitter solution, in the limit of constant Hubble parameter, we first consider the Schwarzschild-de Sitter metric in Painlevé-Gullstrand coordinates Gaur2
$ds^{2}=-(1-\frac{l_{s}}{r}-\frac{r^{2}}{l^{2}})dt^{2}-2\sqrt{\frac{l_{s}}{r}+\frac{r^{2}}{l^{2}}}dtdr+dr^{2}+r^{2}d\Omega_{2}^{2},$ (3)
and factorize the $g_{tt}$ part to its roots as
$1-\frac{l_{s}}{r}-\frac{r^{2}}{l^{2}}=\frac{l_{s}}{l_{1}}(1-\frac{l_{1}}{r})(1-\frac{r}{l_{2}})(1+\frac{r}{l_{1}+l_{2}}).$ (4)
Here $l_{1}$ and $l_{2}$ are respectively the black hole and cosmological horizons which obey
$l_{s}=\frac{l_{1}l_{2}(l_{1}+l_{2})}{(l_{1}+l_{2})^{2}-l_{1}l_{2}},\qquad l^{2}=(l_{1}+l_{2})^{2}-l_{1}l_{2}.$ (5)
Then the the metric is brought to the form
$ds^{2}=-\frac{l_{s}}{l_{1}}\Big{(}1-\frac{l_{1}}{r}-\frac{r^{2}-l_{1}^{2}}{l_{2}(l_{1}+l_{2})}\Big{)}dt^{2}-2\sqrt{\frac{l_{s}}{l_{1}}\frac{l_{1}}{r}+\frac{r^{2}}{l^{2}}}dtdr+dr^{2}+r^{2}d\Omega_{2}^{2}.$ (6)
Relating $l_{2}$ and $l_{s}$ to $l_{1}$ and $l$ we can write this as
$ds^{2}=-\big{(}1-(\frac{l_{1}}{l})^{2}\big{)}\Big{(}1-\frac{l_{1}}{r}-\frac{r^{2}-l_{1}^{2}}{l^{2}-l_{1}^{2}}\Big{)}dt^{2}-2\sqrt{\big{(}1-(\frac{l_{1}}{l})^{2}\big{)}\frac{l_{1}}{r}+\frac{r^{2}}{l^{2}}}dtdr+dr^{2}+r^{2}d\Omega_{2}^{2}.$ (7)
which is useful for our purposes since the dependence on the two horizons is explicit and they are written in a disentangled fashion.
We should note that due to a negative $g_{tr}$ component, this metric in fact represents a white hole in these coordinates Volovik . Using this analogy we propose a metric for a spherically symmetric white hole with a static horizon in cosmology as
$ds^{2}=-(1-\frac{l_{1}^{2}\dot{a}^{2}}{a^{2}})\Big{(}1-\frac{l_{1}}{r}-\frac{r^{2}-l_{1}^{2}}{\frac{a^{2}}{\dot{a}^{2}}-l_{1}^{2}}\Big{)}dt^{2}\\ -2\sqrt{\big{(}1-\frac{l_{1}^{2}\dot{a}^{2}}{a^{2}}\big{)}\frac{l_{1}}{r}+\frac{r^{2}\dot{a}^{2}}{a^{2}}}dtdr+dr^{2}+r^{2}d\Omega_{2}^{2}.$ (8)
The metric (8) can be written in a more suggestive form as
$ds^{2}=-(1-h^{2})dt^{2}-2hdtdr+dr^{2}+r^{2}d\Omega_{2}^{2}.$ (9)
where
$h=\sqrt{\frac{l_{1}}{r}+\frac{r^{3}-l_{1}^{3}}{r}\frac{\dot{a}^{2}}{a^{2}}}.$ (10)
This metric has the property that the white hole and cosmological horizons are detached and it has an static horizon at $r=l_{1}$. Also if we replace $\frac{a}{\dot{a}}$ with de Sitter length $l$ for the cosmological constant, we recover the Schwarzschild-de Sitter metric.
In the large $r$ limit the metric tends to the cosmological metric (2) with $t$ the FLRW time. In the small $r$ limit the metric tends to the Schwarzschild metric (1) with Painlevé-Gullstrand time. We note that using this metric, the FLRW time is compared with Painlevé-Gullstrand time in Schwarzschild metric and not the static Schwarzschild time.
Since the horizon is a little obscure in these coordinates, we can change variables to remove the $g_{tr}$ part of the metric. Writing $t=\gamma(\tau,r)$ the $g_{tr}$ component of the metric becomes
$g_{tr}=-2\dot{\gamma}((1-h^{2})\gamma^{\prime}+h).$ (11)
Therefore to remove this we choose $\gamma$ such that
$\gamma^{\prime}=-\frac{h}{1-h^{2}}.$ (12)
Then the metric becomes
$ds^{2}=-(1-h^{2})\dot{\gamma}^{2}d\tau^{2}+\frac{dr^{2}}{1-h^{2}}+r^{2}d\Omega_{2}^{2}.$ (13)
This metric has a static horizon at $r=l_{1}$ where $h=1$ independent of $\tau$.
We can see that the Ricci scalar and curvature invariants like $R^{\mu\nu}R_{\mu\nu}$ and $R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ remain finite at the horizon. In particular we have at the horizon $r=l_{1}$
$R|_{r=l_{1}}=-6\,l_{1}\frac{\dot{a}^{3}}{a^{3}}+6\,l_{1}\frac{\dot{a}\ddot{a}}{a^{2}}+12\frac{\dot{a}^{2}}{a^{2}}.$ (14)
For visualization we sketch the Ricci scalar, $R^{\mu\nu}R_{\mu\nu}$, $R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ as a function of $r$ for fixed $l_{1}=1$ and $t=10$ for both radiation and matter dominated cases in Figure 1.
Figure 1: The Ricci scalar, $R^{\mu\nu}R_{\mu\nu}$, $R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ as a function of $r$, for $l_{1}=1$ and $t=10$, with blue and orange curves related to radiation and matter respectively. In the last plot, radiation and matter curves are almost coincident. These plots represent the observation that this metric has a singularity only at $r=0$ and not at the event horizon.
Another interesting property is that $T^{t}_{r}$ which we find from requiring Einstein's equation is zero. In other words there is no flux of energy in radial direction. This is the requirement also present in McVittie's solution. On the other hand, the large $r$ limit of stress tensor tends to the cosmological values.
Explicitly we have
$G^{t}_{t}=-3\frac{\dot{a}^{2}}{a^{2}},\quad\quad G^{t}_{r}=0,$ (15)
$G^{r}_{t}=-2\,r\big{(}1-\frac{l_{1}^{3}}{r^{3}}\big{)}\Big{(}\frac{\dot{a}^{3}}{a^{3}}-\frac{\dot{a}\ddot{a}}{a^{2}}\Big{)}=g^{tr}(G^{r}_{r}-G^{t}_{t}),$ (16)
and
$G^{\theta}_{\theta}=G^{\phi}_{\phi}=-3\frac{\dot{a}^{2}}{a^{2}}-\frac{l_{1}(l_{1}^{3}-7r^{3})}{2r^{3}g_{tr}^{3}}\Big{(}\frac{\ddot{a}\dot{a}}{a^{2}}-\frac{\dot{a}^{3}}{a^{3}}\Big{)}-\frac{(l_{1}^{3}-r^{3})(l_{1}^{3}-4r^{3})}{2r^{3}g_{tr}^{3}}\Big{(}\frac{\dot{a}^{5}}{a^{5}}-\frac{\ddot{a}\dot{a}^{3}}{a^{4}}\Big{)}$ (17)
We can see that (assuming the gravitational theory is general relativity,) the radial pressure for small $r$ is not equal to angular pressure, i.e. we do not have $p_{\perp}=p_{r}$. This shows that the metric close to the black hole is not isotropic at every point. But since the $G^{\mu}_{\nu}$ tends to the cosmological value for large $\frac{r}{l_{1}}$ this property ($p_{\perp}=p_{r}$) is restored in this limit.
2.2 A systematic derivation for the white hole case
Having found the metric (8), we show that it could have been derived in a more systematic way uniquely. We begin with the Painlevé-Gullstrand-like metric ansatz
$ds^{2}=-f(t,r)dt^{2}-2h(t,r)dtdr+dr^{2}+r^{2}d\Omega_{2}^{2},$ (18)
and consider the following assumptions:
$a)$ There is no flux of energy in radial direction, i.e. we assume $T^{t}_{r}=0$.
$b)$ There is a static horizon at $r=l_{1}$, such that $g^{rr}|_{r=l_{1}}=0$.
$c)$ We assume the energy density is the same as the FLRW one, i.e. we assume $G^{t}_{t}=-3\frac{\dot{a}^{2}}{a^{2}}$.
$d)$ The metric at large $r$ tends to the cosmological FLRW metric.
We note that compared to McVittie's assumptions, we have dropped the $p_{\perp}=p_{r}$ assumption and instead we have added assumption b.
Using the metric (18), we have
$G^{t}_{r}=-\frac{h(f^{\prime}+2hh^{\prime})}{r(f+h^{2})^{2}},$ (19)
where denotes derivative with respect to $r$. Requiring this to vanish considering the first assumption, we find
$f(t,r)=\gamma(t)-\omega(t,r),$ (20)
where we have defined $\omega(t,r)=h^{2}(t,r)$. We then consider the second assumption. From this and using (20) we require
$\omega(t,l_{1})=\gamma(t).$ (21)
Considering assumption c, we have
$G^{t}_{t}=-\frac{\omega(t,r)+r\omega^{\prime}(t,r)}{r^{2}\omega(t,l_{1})}=-3\frac{\dot{a}^{2}}{a^{2}}.$ (22)
Integrating we find
$\omega(t,r)=\omega(t,l_{1})r^{2}\frac{\dot{a}^{2}}{a^{2}}+\frac{c(t)}{r}.$ (23)
Considering assumption d we have $\omega(t,l_{1})=1$ and from this considering (23) at $r=l_{1}$, we have
$c(t)=l_{1}\big{(}1-\frac{\dot{a}^{2}}{a^{2}}l_{1}^{2}\big{)}.$ (24)
Therefore with these assumptions the proposed metric (8) is unique.
2.3 Energy conditions
In this section we study the energy conditions for the stress tensor corresponding to this metric assuming the gravitational theory is general relativity.
We begin with the weak energy condition that states for every timeline vector $k^{\mu}$ we should have
$T_{\mu\nu}k^{\mu}k^{\nu}\geq 0$ (25)
Considering the case where only $k^{t}=1$ and $k^{r}$ are nonzero we have for the vector $k^{\mu}$ the inequality
$g_{tt}+2g_{tr}k^{r}+g_{rr}(k^{r})^{2}<0$ (26)
which using $g_{tt}=-1+g_{tr}^{2}$ leads to
$-1-g_{tr}<k^{r}<1-g_{tr}.$ (27)
Using these marginal values of $k^{r}$, we find that in both cases
$T_{\mu\nu}k^{\mu}k^{\nu}=\frac{2(l_{1}^{3}-r^{3})(\dot{a}^{2}-a\ddot{a})}{r^{2}a^{3}g_{tr}}.$ (28)
For a typical power-law scale factor, this is always positive outside the horizon and negative inside the horizon (since $g_{tr}<0$). Therefore assuming the gravitational theory is general relativity, the stress tensor corresponding to this metric does not satisfy the weak and null energy conditions inside the event horizon.
Considering the strong energy condition we find that it is violated in the vicinity of the white hole (see figure 2).
Figure 2: Here we have plotted $T_{tt}-\frac{1}{2}g_{tt}T$ as a function of $r$. We have chosen $l_{1}=1$ and $t=10$. The blue and orange curves respectively correspond to radiation and matter dominated scale factors. The strong energy condition is violated in the white hole's vicinity.
2.4 NEC and static horizon
Here we drop assumption c mentioned in the previous section to see if we can find a metric with a static horizon that respects NEC. We begin with the metric
$ds^{2}=-f(t,r)dt^{2}-2h(t,r)dtdr+dr^{2}+r^{2}d\Omega_{2}^{2}.$ (29)
We still have (20), which considering assumption d takes the form
$f(t,r)=1-h^{2}(t,r).$ (30)
We then consider NEC. It can be shown that NEC requires $\dot{h}(t,r)<0$ everywhere, (as we have $T_{\mu\nu}k^{\mu}k^{\nu}=-2\frac{\dot{h}}{r}$). Since we assume to have a static horizon at some $r=l_{1}$ (for which $h(t,l_{1})=1$), at the horizon $\dot{h}$ becomes zero and the only way it does not become positive below the horizon is to have $\dot{h}^{\prime}(t,l_{1})=0$ at the horizon.
One possible metric which respects NEC and assumptions a,b,d has
$h(t,r)=\sqrt{\frac{l_{1}}{r}+r^{2}(1-\frac{l_{1}^{3}}{r^{3}})\frac{\dot{a}^{2}}{a^{2}}+l_{1}^{2}(1-\frac{l_{1}^{3}}{r^{3}})(\frac{1}{l_{0}^{2}}-\frac{\dot{a}^{2}}{a^{2}})}.$ (31)
Here $l_{0}$ is related to the cosmological constant $\Lambda=\frac{3}{l_{0}^{2}}$.
We have
$\dot{h}=-\frac{(l_{1}-r)^{2}(l_{1}+r)(l_{1}^{2}+l_{1}r+r^{2})}{r^{3}h}\big{(}\frac{\dot{a}^{3}}{a^{3}}-\frac{\dot{a}\ddot{a}}{a^{2}}\big{)},$ (32)
which is negative for all $r$.
2.5 The feasibility of a static horizon, the black hole case
In this section, we consider a more generic form of the metric that can represent a black hole in cosmology with a static horizon. For this, we consider the metric ansatz Volovik
$ds^{2}=-\big{(}N^{2}-g_{rr}h^{2}\big{)}dt^{2}+g_{rr}dr^{2}+2g_{rr}h\,dtdr+r^{2}d\Omega_{2}^{2},$ (33)
We assume $N^{2}=\frac{1}{g_{rr}}=\lambda(t,r)$.
Considering the assumption that $G^{t}_{t}=-3H^{2}$, which is its cosmological value, we can write
$\partial_{r}(\frac{rh^{2}}{\lambda}+r-r\lambda)=H^{2}\partial_{r}(r^{3}),$ (34)
$h^{2}=\lambda(\lambda-1+H^{2}r^{2}+\frac{\alpha(t)}{r})$ (35)
Using the requirement that we have a static horizon at $r=l_{1}$, we find
$g^{rr}=\lambda(r)-\frac{h^{2}}{\lambda}|_{r=l_{1}}=0.$ (36)
From this we have $\alpha(t)=l_{1}(1-H^{2}l_{1}^{2})$ and
$h^{2}=\lambda\big{(}\lambda-1+H^{2}r^{2}+\frac{l_{1}}{r}(1-H^{2}l_{1}^{2})\big{)}.$ (37)
In accordance with Volovik , we can choose
$\lambda=1-3(\frac{l_{s}H}{2})^{2/3},\qquad l_{s}=l_{1}(1-H^{2}l_{1}^{2}).$ (38)
Then we have
$h^{2}=(1-\lambda)\lambda(r-r_{0})^{2}\frac{r+2r_{0}}{3\,r\,r_{0}^{2}},\qquad r_{0}=(\frac{l_{s}}{2H^{2}})^{1/3}.$ (39)
Therefore we can choose different signs of the square root for $r>r_{0}$ and $r<r_{0}$.
Again one can show that this result violates NEC. Similar to the white hole case we can drop the assumption that the energy density is equal to the cosmological value and find a metric that also satisfies NEC.
3 The case of constant equation of state
In this section we consider the possibility that the metric respects the cosmological stress tensor equation of state.
3.1 The white hole case
Here we use the metric ansatz (18). Assuming $T^{t}_{r}=0$, from (19) we find
$f(t,r)=\gamma(t)-h^{2}(t,r).$ (40)
We do not consider any further assumptions and continue our analysis for the case of constant equation of state.
We assume $p_{\perp}=p_{r}=\omega\rho$. Assuming $G^{r}_{t}=g^{tr}(G^{r}_{r}-G^{t}_{t})$  [^1] This follows from $G^{t}_{r}=0$ and $G^{tr}=G^{rt}$., we have from Einstein equations
$G^{r}_{t}=\frac{h}{r\gamma^{2}}(2\gamma\dot{h}-h\dot{\gamma})=-\frac{h}{\gamma^{2}}\kappa_{4}(1+\omega)\rho,$ (41)
and therefore
$\frac{1}{r}(2\gamma\dot{h}-h\dot{\gamma})=-\kappa_{4}(1+\omega)\rho.$ (42)
On the other hand from $G^{t}_{t}=-\kappa_{4}\rho$ and (42), we find
$h^{2}+2rhh^{\prime}+\frac{2r\gamma^{2}\dot{h}-rh\gamma\dot{\gamma}}{1+\omega}=0.$ (43)
We also have by considering $G^{r}_{r}=\kappa_{4}\omega\rho$,
$h^{2}+2rhh^{\prime}+\frac{2r\dot{h}-rh\frac{\dot{\gamma}}{\gamma}}{1+\omega}=0.$ (44)
Using (43) and (44) we find
$(2r\dot{h}-rh\frac{\dot{\gamma}}{\gamma})(\gamma^{2}-1)=0.$ (45)
So we either have $\gamma^{2}=1$ or $2\frac{\dot{h}}{h}=\frac{\dot{\gamma}}{\gamma}$.
The second possibility requires $h^{2}=c(r)\gamma(t)$. Using (43) again we find $c(r)=\frac{l_{s}}{r}$ where $l_{s}$ is a constant length and the metric takes the form
$ds^{2}=-\gamma(t)(1-\frac{l_{s}}{r})dt^{2}-2\sqrt{\frac{l_{s}}{r}\gamma(t)}dtdr+dr^{2}+r^{2}d\Omega^{2},$ (46)
which is the Schwarzschild white hole metric in Painlevé-Gullstrand coordinates (This is evident by changing $t$ to $t^{\prime}$ such that $dt^{\prime}=\sqrt{\gamma(t)}dt$). Therefore, in order to find a cosmological white hole metric, our preference will be $\gamma(t)=1$ in what follows.
Assuming $\gamma(t)=1$, we have from the $G^{0}_{0}$ component of Einstein equations and (42)
$h^{2}+2rhh^{\prime}+\frac{2}{1+\omega}r\dot{h}=0.$ (47)
Considering the $G^{\theta}_{\theta}$ component we find
$\partial_{r}(r\dot{h}+rhh^{\prime}+\frac{h^{2}}{2})=-\kappa_{4}r\omega\rho=\frac{1}{r}(h^{2}+2rhh^{\prime}+2r\dot{h}),$ (48)
where we used the $G^{r}_{r}$ component as well. In the following, we consider the case of $\omega=0$ and $\omega\neq 0$ separately.
3.1.1 The case of $\omega\neq 0$
In this case, from (48) we find
$h^{2}+2rhh^{\prime}+2r\dot{h}=\frac{2\omega}{1+\omega}\dot{\zeta}(t)r^{2}.$ (49)
Here $\zeta(t)$ is some function of $t$ and we have chosen the coefficient and a derivative with respect to $t$ for later convenience.
Subtracting (47) from (49), we further find
$h=\zeta(t)r+\alpha(r).$ (50)
Using (47) again we find
$3\zeta^{2}(t)r^{2}+4r\zeta(t)\alpha(r)+2r^{2}\zeta(t)\alpha^{\prime}(r)+\frac{2}{1+\omega}r^{2}\dot{\zeta}+\alpha^{2}(r)\\ +2r\alpha(r)\alpha^{\prime}(r)=0$ (51)
The only nontrivial solution of this equation which is of cosmological interest is $\alpha(r)=0$, $\zeta(t)=\frac{2}{3(1+\omega)t}$. Therefore we do not have a cosmological solution that satisfies $p_{r}=p_{\perp}=\omega\rho$ for $\omega\neq 0$.
3.1.2 The case of matter with $\omega=0$
To simplify our analysis, we change the coordinates to a semi-homogeneous form by setting $r=a(t,R)R$, demanding the transformed metric to be diagonal by setting $h=R\dot{a}$. Therefore, the metric becomes simply
$ds^{2}=-dt^{2}+(a+Ra^{\prime})^{2}dR^{2}+a^{2}R^{2}d\Omega_{2}^{2},$ (52)
where $a^{\prime}=\partial_{R}a$. We note that if $a$ was independent of $R$, this was the FLRW metric. In this coordinate $T^{R}_{t}=T^{t}_{R}=0$ and the Einstein equations become
$\frac{3a\dot{a}^{2}+R\,a^{\prime}\dot{a}^{2}+2R\,a\dot{a}\dot{a}^{\prime}}{a^{2}(a+R\,a^{\prime})}=\frac{\partial_{R}(R^{3}a\dot{a}^{2})}{R^{2}a^{2}(a+R\,a^{\prime})}=\kappa_{4}\rho,$ (53)
$(\frac{\dot{a}}{a})^{2}+2\frac{\ddot{a}}{a}=-\kappa_{4}p_{r},$ (54)
$\frac{\dot{a}^{2}+R\,\dot{a}\dot{a}^{\prime}+2a\ddot{a}+R\,a^{\prime}\ddot{a}+R\,a\ddot{a}^{\prime}}{a(a+R\,a^{\prime})}=-\kappa_{4}p_{\perp}.$ (55)
In this section we consider the case of a Schwarzschild white hole immersed in cosmological matter. In this case we have $p_{r}=p_{\perp}=0$. From stress energy conservation we can find the structure of $\rho$. We have
$\frac{\dot{\rho}}{\rho}+2\frac{\dot{a}}{a}+\frac{\dot{a}+R\,\dot{a}^{\prime}}{a+R\,a^{\prime}}=0,$ (56)
and we find
$\kappa_{4}\rho=\frac{\beta(R)}{a^{2}(a+R\,a^{\prime})}.$ (57)
From (53) we can write
[…]
From (54) we find
$a(t,R)=c(R)\,\big{(}\frac{3}{2}t-b(R)\big{)}^{2/3},$ (59)
which also satisfies (55). Using (58) this means
[…]
We first re-derive the Schwarzschild metric in these coordinates. If we only had a point mass contribution to $\rho$, this equation would lead to
$c(R)=\frac{l_{s}^{1/3}}{R}$ (61)
or
$a(t,R)=\frac{l_{s}^{1/3}}{R}(\frac{3}{2}t-b(R))^{2/3}.$ (62)
In our original central coordinate this means that we had
$h=R\,\dot{a}=\frac{l_{s}^{1/3}}{(\frac{3}{2}t-b(R))^{1/3}}=\sqrt{\frac{l_{s}}{r}}$ (63)
where in the last equality we used $r=R\,a$. We see that this simple case corresponds to the Schwarzschild metric in Painlevé-Gullstrand coordinates. We also note that in this coordinate the singularity is at the surface $\frac{3}{2}t-b(R)=0$. There is a freedom in our choice of $b(R)$.
Now returning to the case of matter-dominated cosmology we can add a term related to the cosmological matter density to $\beta(R)$. We note that in the asymptotic case we should have $\kappa_{4}\beta(R)=\beta_{0}$ where $\beta_{0}$ is a constant. Considering that this is also the case for large $t$ we set $\kappa_{4}\beta(R)=\beta_{0}$ and find
$c(R)=\left(\frac{l_{s}}{R^{3}}+\frac{\beta_{0}}{3}\right)^{1/3},$ (64)
and
$h=R\dot{a}=\frac{\left(l_{s}+\frac{\beta_{0}}{3}R^{3}\right)^{1/3}}{\left(\frac{3}{2}t-b(R)\right)^{1/3}}=\frac{r}{\frac{3}{2}t-b(R)}.$ (65)
The apparent horizon is at $h=1$. Therefore we have
$r_{H}=\frac{3}{2}t-b(R_{H}).$ (66)
On the other hand, using $r=Ra$ we have
$r_{H}=(l_{s}+\frac{\beta_{0}}{3}R^{3})^{1/3}(\frac{3}{2}t-b(R))^{2/3},$ (67)
and using (66) we can write
$r_{H}=l_{s}+\frac{\beta_{0}}{3}R_{H}^{3}.$ (68)
As an example, if we choose $b(R)=\frac{\kappa l_{s}^{4}}{R^{3}}-d$, using (66) and (68) we find
$r_{H}=l_{s}+\frac{\frac{\beta_{0}}{3}\kappa l_{s}^{4}}{\frac{3}{2}t+d-r_{H}},$ (69)
with solutions
$r_{H}(t)=\frac{1}{4}\big{(}3t+2d+2l_{s}\pm\sqrt{(3t+2d-2l_{s})^{2}-\frac{16}{3}\beta_{0}\kappa l_{s}^{4}}\big{)}.$ (70)
For large $t$ one of these tends to $r_{H}=\frac{3t}{2}$, which is the cosmological horizon, and the other tends to $r_{H}=l_{s}$, which is the white hole horizon. Therefore, for large $t$ we effectively have a static white hole horizon.
We had freedom in our choice of $b(R)$. Requiring that the metric tends to the cosmological metric for large $R$ will force us to choose $b(R)$ such that it tends to a constant for large $R$. The surface $\frac{3t}{2}=b(R)$, $r=0$ is where both the cosmological singularity and white hole singularity reside.
We can prove that generically the resulting apparent horizon in this case will be dynamical (apart from the asymptotic behavior). To see this, we write
$h=R\dot{a}=\frac{Rc(R)}{(\frac{3t}{2}-b(R))^{1/3}}=\frac{r}{\frac{3t}{2}-b(R)}.$ (71)
Requiring $h=1$ at the horizon and considering $r=Ra$, we can write
[…]
$r_{H}=\frac{3t}{2}-b(R_{H}).$ (73)
Taking a time derivative, we find
$\dot{r}_{H}=\dot{R}_{H}R_{H}^{2}\beta(R_{H})=\frac{3}{2}-\dot{R}_{H}b^{\prime}(R_{H}).$ (74)
From this, we can see that it is impossible to have $\dot{r}_{H}=0$ exactly. [^2] To have $\dot{r}_{H}=0$ from the first equality we should either have $\dot{R}_{H}=0$ which contradicts the second equality or we should have $\beta(R_{H})=0$ which again leads to $R_{H}=const.$ which is a contradiction.
3.1.3 The case of matter plus cosmological constant
In this case from $-\kappa_{4}p_{r}=\frac{3}{l^{2}}$ and (54) we find
$a(t,R)=c(R)\sinh^{2/3}\left(\frac{3t}{2l}-\frac{b(R)}{l}\right).$ (75)
We again have (56) and (57) for the matter part of energy density. From (53) we can write
$\partial_{R}(R^{3}a\dot{a}^{2})=R^{2}\beta(R)+\frac{3}{l^{2}}R^{2}a^{2}(a+Ra^{\prime}).$ (76)
Using (75) we find
$\frac{3}{l^{2}}c^{2}(R)(c(R)+Rc^{\prime}(R))=\beta(R)\,,$ (77)
or alternatively
$\frac{1}{l^{2}}\partial_{R}(R^{3}c^{3})=R^{2}\beta(R).$ (78)
We are again left with two unfixed functions of $R$, namely $b(R)$ and $\beta(R)$. One can use the degree of freedom in choosing the $R$ coordinate to fix one of them.
If we assume $\beta(R)=\beta_{0}$ to be a constant we find
$c(R)=\left(\frac{l_{s}l^{2}}{R^{3}}+\frac{\beta_{0}\,l^{2}}{3}\right)^{1/3}$ (79)
By studying $h=R\,\dot{a}$ one can see that for $\beta_{0}=0$, which is the case of pure cosmological constant, $l_{s}$ is the Schwarzschild radius, i.e. for $\beta_{0}=0$ we have
$h|_{\beta_{0}=0}=R\,\dot{a}=\sqrt{\frac{l_{s}}{r}+\frac{r^{2}}{l^{2}}}.$ (80)
We study the apparent horizon, which is located at $h=1$. We have
$h=R\,\dot{a}=\frac{r}{l}\coth\left(\frac{3t}{2l}-\frac{b(R)}{l}\right).$ (81)
From this at $h=1$ we find
$\sinh^{2}\left(\frac{3t}{2l}-\frac{b(R)}{l}\right)=\frac{\frac{r_{H}^{2}}{l^{2}}}{1-\frac{r_{H}^{2}}{l^{2}}}.$ (82)
Using (75) and (79) and assuming $b(R)$ behaves such that the location of the horizon for large $t$ goes to small $R_{H}$, we can write for large $t$
$r_{H}\approx(l_{s}l^{2})^{1/3}\Big{(}\frac{\frac{r_{H}^{2}}{l^{2}}}{1-\frac{r_{H}^{2}}{l^{2}}}\Big{)}^{1/3},$ (83)
and
$\frac{l_{s}}{r_{H}}+\frac{r_{H}^{2}}{l^{2}}=1.$ (84)
The smallest root of this equation is the white hole horizon. This means that the location of the white hole horizon for large $t$ tends to its Schwarzschild-de Sitter counterpart.
If we require that the observer is at large $R$ then we should choose $b(R)$ such that the cosmological behavior is dominant for large $R$. The location of singularity is at $\frac{3t}{2}=b(R)$. Choosing as an example $b(R)=\frac{\kappa l_{s}^{4}}{R^{3}}-d$, the location of singularity tends to $R=0$ for large $t$ and to a constant $R$ for small $t$.
3.2 The Black Hole case
In this section we study the case of black hole immersed in cosmological matter and matter plus cosmological constant.
3.2.1 The case of matter
Beginning from the metric ansatz
$ds^{2}=-N^{2}dt^{2}+g_{rr}(dr-hdt)^{2}+r^{2}d\Omega_{2}^{2},$ (85)
we find after a change of variable $r=f(t,R)$ such that $h=\dot{f}$ the metric
$ds^{2}=-N^{2}dt^{2}+g_{rr}(f^{\prime})^{2}dR^{2}+f^{2}d\Omega_{2}^{2}.$ (86)
Assuming we only have a dust contribution to stress tensor such that only $T^{t}_{\,t}=-\kappa_{4}\rho$ is nonzero, we find from stress tensor conservation
$N^{\prime}=0.$ (87)
Therefore $N$ is only a function of $t$ and by redefining $t$ we can set it to $N=1$.
Requiring $T^{t}_{\,R}=0$ we further find $\dot{g}_{rr}=0$ and we define $g_{rr}=\frac{1}{\lambda^{2}(R)}$. From $T^{R}_{\,R}=0$ we find
$\partial_{t}\big{(}(1-\lambda^{2})f+f\dot{f}^{2}\big{)}=0,$ (88)
On the other hand from conservation equation we can fix the form of $\kappa_{4}\rho$ to be
$\kappa_{4}\rho=\frac{\beta(R)}{f^{2}f^{\prime}}$ (89)
and from $T^{t}_{\,t}=-\kappa_{4}\rho$ we find
$\partial_{R}\big{(}(1-\lambda^{2})f+f\dot{f}^{2}\big{)}=\beta(R).$ (90)
Therefore we can write this as
$f\dot{f}^{2}=\zeta(R)-\omega(R)f$ (91)
where we have defined $\zeta(R)$ such that $\zeta^{\prime}(R)=\beta(R)$ and $\omega(R)=1-\lambda^{2}(R)$.
The general solution to this equation can be written as
$\frac{\zeta}{\omega\sqrt{\omega}}\arcsin{\sqrt{\frac{\omega f}{\zeta}}}-\frac{\sqrt{f(\zeta-\omega f)}}{\omega}=\pm t+b_{\pm}(R).$ (92)
We note that (91) requires $f<\frac{\zeta}{\omega}$ for $\omega>0$.
We study the behavior of this solution in different limits. For $f\ll\frac{\zeta}{\omega}$ the solution becomes
$f=\zeta^{1/3}(R)(\frac{3}{2}(\pm t+b_{\pm}(R)))^{2/3}.$ (93)
In the limit where $f\approx\frac{\zeta}{\omega}$ the first trem in (92) is dominant and the solution takes the form
$f=\frac{\zeta}{\omega}\sin^{2}\big{(}\frac{\omega^{3/2}}{\zeta}(\pm t+b_{\pm}(R))\big{)},$ (94)
such that at ±$\pm t+b_{\pm}(R)=\frac{\pi\,\zeta(R)}{2\,\omega^{3/2}(R)}$, $f$ reaches its maximum for a given $R$.
To understand the behavior of the solution we define $f=\frac{\zeta^{1/3}}{d}\hat{f}$ where $d$ has dimension length and assume $\omega(R)=(\frac{\zeta}{d})^{2/3}$ to find
$\hat{f}\dot{\hat{f}}^{2}=d-\hat{f},$ (95)
with solutions
$d\arcsin{\sqrt{\frac{\hat{f}}{d}}}-\sqrt{\hat{f}(d-\hat{f})}=\pm t+b_{\pm}(R).$ (96)
For a fixed $R$, $b_{\pm}(R)$ are constants. If we set $b_{\pm}=0$, then the solutions are as in figure 3.
[…]
To have a reasonable solution that moves from one solution smoothly to the other, we note that $\hat{f}$ takes its maximum value at ±$\pm t^{c}_{\pm}+b_{\pm}=\frac{\pi\,d}{2}$ where $t^{c}_{\pm}$ are transition times. Therefore for a smooth transition we need to have $t^{c}_{-}=t^{c}_{+}=t^{c}$ from which we find $b_{+}+b_{-}=\pi\,d$. For $b_{+}=0$ and $b_{-}=\pi\,d$ the plot becomes as in figure 4.
Figure 4: The plot for $\hat{f}$ with $b_{\pm}$ chosen such that we have a smooth transition between the $+$ and -- branches.
According to this, the behavior of $\hat{f}$ is such that at each value of $R$, for early time ($t<t^{c}=\frac{\pi\,\zeta}{2\,\omega^{3/2}}$) we have an increasing scale factor reminiscent of a an expanding cosmology. After the crossover time the scale factor becomes decreasing which is reminiscent of the behavior in the vicinity of a black hole.
According to (93) the metric away from the transition point only depends on $\zeta$ and $b_{\pm}$. Therefore our calculations regarding the apparent horizon remain the same as the white-hole case and we can have an apparent horizon which tends to a constant for large time.
3.2.2 The case of matter plus cosmological constant
Similar to the previous section we can again set $N=1$ and from $T^{t}_{R}=0$ we find $\dot{\lambda}=0$. From $\kappa_{4}T^{R}_{R}=-\frac{3}{l_{0}^{2}}$ related to the cosmological constant we can write
$\partial_{t}\big{(}(1-\lambda^{2}(R))f+f\dot{f}^{2}-\frac{f^{3}}{l_{0}^{2}}\big{)}=0.$ (97)
On the other hand from energy momentum conservation we find as before
$\kappa_{4}\rho=\frac{\beta(R)}{f^{2}f^{\prime}}.$ (98)
From $\kappa_{4}T^{t}_{t}=-\kappa_{4}\rho-\frac{3}{l_{0}^{2}}$ we find
$(1-\lambda^{2})f+f\dot{f}^{2}-\frac{f^{3}}{l_{0}^{2}}=\zeta(R),$ (99)
where we have $\zeta^{\prime}(R)=\beta(R)$.
We study this equation in two limits. Close to the black hole we expect from previous results that $f\ll\frac{\zeta}{\omega}$ (defining $\omega=1-\lambda^{2}$). In this limit we have
$f\dot{f}^{2}-\frac{f^{3}}{l_{0}^{2}}\approx\zeta(R),$ (100)
and we find
$f=(l_{0}^{2}\zeta)^{1/3}\sinh^{2/3}\big{(}\pm\frac{3t}{2l_{0}}+\frac{b(R)}{l_{0}}\big{)}.$ (101)
Another important limit is where $f\approx\frac{\zeta}{\omega}$. This is close to the region where the sign of $h=\dot{f}$ changes. In order to understand the behavior of $f$ in this region, we find it suitable to set define $f=\big{(}\frac{\zeta}{d}\big{)}^{1/3}\hat{f}$ and choose $\omega=\big{(}\frac{\zeta}{d}\big{)}^{2/3}$ to find
$\hat{f}+\hat{f}\dot{\hat{f}}^{2}-\frac{\hat{f}^{3}}{l_{0}^{2}}=d.$ (102)
To study the region close to $f=\frac{\zeta}{\omega}$ we write $\hat{f}=d+\phi$ where $\phi\ll d$ and find (we assume $d\ll l_{0}$)
$\phi+d\,\dot{\phi}^{2}\approx\frac{d^{3}}{l_{0}^{2}},$ (103)
with solution
$\phi=\frac{d^{3}}{l_{0}^{2}}-\frac{d}{4}\big{(}\frac{t}{d}-l_{0}c(R)\big{)}^{2}.$ (104)
Here $c(R)$ is an arbitrary function of $R$. We see that this has the same behavior as in figure 4 close to the cross over point $t_{c}=l_{0}\,d\,c(R)$.
4 Summary
Despite the recent claim that black hole horizons should be cosmologically coupled, in this work, we provide counterexamples for cosmological black hole metrics that have static or asymptotically static horizons.
We find that it is technically much easier to find white hole metrics in an expanding cosmology or a black hole metric in a contracting cosmology. This is due to the fact that the sign of $h$ in (18) is negative for a black hole and positive for an expanding cosmology. Therefore, to find a black hole metric embedded in expanding cosmology, we need to have a sign change in $h$ in the vicinity of the black hole relative to the asymptotic cosmology.
Due to this complexity, we first studied the more trivial case of a white hole in expanding cosmology and then dealt with the issue of the sign change in $h$ by generalizing the form of the metric. The generalization is effectively a curvature-like behavior (relative to the flat FLRW metric).
In the first part of this work, we propose white and black hole metrics with static horizons embedded in FLRW cosmology. These metrics have the following properties:
$a)$ It has a static horizon with no curvature singularity.
$b)$ It tends to the cosmological FLRW metric for large physical radius $r$.
$c)$ The energy density is the same as the cosmological case, but the pressure is different.
$d)$ The stress tensor also tends to the FLRW values for large $r$.
$e)$ It violates NEC. But we show that it is in principal possible to drop assumption c and find metrics with static horizons that satisfy NEC as well.
The corresponding stress tensor, assuming general relativity, violates the strong energy condition in the vicinity of the black hole.
$f)$ For the case of a constant Hubble parameter, it is equal to the Schwarzschild-de Sitter metric.
In the second part of this work we considered a more realistic framework with a constant equation of state (for the cases of matter and radiation) and first study the white hole case due to its simplicity. We show that we have no non-trivial solutions for the case of $\omega\neq 0$. For the case of matter with $\omega=0$, we derived the white hole metric consistent with this assumption. The crucial feature of this metric is that, despite having an asymptotically vanishing Hubble parameter, we have no singularities at the horizon.
We also studied the case of matter plus cosmological constant, where we assumed the pressure component remains equal to the constant cosmological value. In all of these cases, we show that the white hole horizon can asymptotically (for large $t$) tend to a static horizon.
Finally we considered the case of black hole immersed in cosmological matter (with and without cosmological constant) with a velocity function $h$ that changes sign in the vicinity of the black hole.
Acknowledgment
We thank Mohammad Mehdi Sheikh-Jabbari for comments and discussions. We specially thank Grigorii E. Volovik for mentioning his important work that helped us improve this manuscript.